How do you find the answer: two numbers added together equal 56 and the difference of the same two numbers is 8. What is the best way to find this answer
Answer by josgarithmetic(39620)

The best way to answer is through simple algebra. A bad way to answer is trial & error.

First translate into variables and expressions. two numbers, x, and y. Sum is 56, difference is 8.

x+y=56, and y-x=8, or x-y=8 depending on which you want as greater.

from y-x=8, add x to both sides getting y=8+x.
Substitute into the sum equation and solve for x.
x+(8+x)=56
x+8+x=6=56
2x=56-8
2x=48, x=6, so what is y?